OpenStudy (anonymous):

find the negative value of x where the tangent line to the graph of f(x)=3*x+(1/x) has a slope=-2?

zepdrix (zepdrix):

So we want to find an x value for which f'(x)=-2. They said `find the negative value of x`, which implies we'll probably find more than one x value when we get to our solution.

zepdrix (zepdrix):

Try taking the derivative, what do you get? :) f'(x)=?

OpenStudy (anonymous):

well 3*x is just 3, and 1/x is -1/x^2

zepdrix (zepdrix):

\[\Large f'(x)\quad=\quad 3-\frac{1}{x^2}\] Ok good, so we want to find a location where the slope of the tangent line is -2.\[\Large -2\quad=\quad3-\frac{1}{x^2}\]

OpenStudy (anonymous):

ok i tried this earlier but i must have just done something wrong. so do you just solve for x?

zepdrix (zepdrix):

Yes, solve for x. See how we have a square on the x? We'll end up getting 2 solutions, positive and negative. We'll want to take the negative one as our final answer.

OpenStudy (anonymous):

so can't you bring the x^2 into the numerator just by making it negative?

zepdrix (zepdrix):

There are a bunch of different ways to solve for x, just depends what you are most comfortable with. Personally I would multiply both sides by x^2 giving us,\[\Large -2x^2\quad=\quad 3x^2-1\] But that's cause I don't like dealing with those ugly fractions.

zepdrix (zepdrix):

Yes you could change x^2 to x^{-2} and bring it to the numerator, but I don't think that's the route you wanna take :p

OpenStudy (anonymous):

so how did you get the x^2 on the -2 as well?

zepdrix (zepdrix):

Multiplying both sides by x^2 gives us,\[\Large x^2\left(-2\right) \quad=\quad x^2\left(3-\frac{1}{x^2}\right)\] Which simplifies to,\[\Large -2x^2\quad=\quad 3x^2-1\]

OpenStudy (anonymous):

sorry it took me so long to respond, the website stopped working

zepdrix (zepdrix):

yes it did :( grumble grumble..

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

so can we combine -2*x^2 and 3*x^2?

zepdrix (zepdrix):

mhm

zepdrix (zepdrix):

\[\Large 0=5x^2-1\]

OpenStudy (anonymous):

so its + or - 1/5?

zepdrix (zepdrix):

Woops! Don't forget to take the square root! :O Maybe just a typo. \[\Large x\quad=\quad\pm \frac{1}{\sqrt5}\quad=\quad\pm\frac{\sqrt5}{5}\]

zepdrix (zepdrix):

We should probably write it like that ^ to be legit.

zepdrix (zepdrix):

So which value of x do we want? The positive or negative? +_+

OpenStudy (anonymous):

ok i got it and checked it and its all good!!! the - one of course!

zepdrix (zepdrix):

yay team \c:/

OpenStudy (anonymous):

yay! thank you!!! now i can sleep haha